Two 4-sided dice are tossed 10 times and the sum of the “down” faces is noted. What is the probability that the sum of 4 occurs at most two times? Leave your answer as a percent rounded to one decimal place. Provide evidence of your work by typing out your full solution.


1
Expert's answer
2021-03-17T03:11:27-0400

Solution:

Let "X" be the random variable denoting a sum of 4 on two 4-sided dice.

Total number of possible outcomes"=4^2=16"

Favorable outcomes"=\{(1,3),(3,1),(2,2)\}"

Total number of favorable outcomes"=3"

"p=\dfrac 3{16}, q=1-\dfrac 3{16}=\dfrac {13}{16},n=10"

"X\sim \text{Binomial}(10,\dfrac 3{16})"

So, "P(X\le2)=P(X=0)+P(X=1)+P(X=2)"

"=^{10}C_0(\dfrac 3{16})^0(\dfrac {13}{16})^{10}+^{10}C_1(\dfrac 3{16})^1(\dfrac {13}{16})^{9}+^{10}C_2(\dfrac 3{16})^2(\dfrac {13}{16})^{8}"

"\approx0.7152=71.52\%"